The sine wave represents Voltage Eab across the terminal a, b of an AC motor that operates at 50 Hz. Knowing that 0=30 and Em = 100V, find the voltage at t=0 and t= 27.144 S.


1
Expert's answer
2021-06-21T16:26:25-0400

As per the given question,



The voltage at t= 0 is,

"E_{ab}=E_m \cos(360ft+\theta)"

Now, substituting the values,

"=100\cos(360\times 50\times0+30^\circ)" V

"=100\cos(30^\circ)" V

"=86.6V"

The voltage at t= 27.144

"\Rightarrow E_{ab}=100\cos(360\times 50\times 27.144+30^\circ)"

"\Rightarrow E_{ab}=100\cos(488622^\circ)"

"\Rightarrow E_{ab}=-20.8V"